The earth revolves round the sun with constant speed on an orbit of radius r=1.5x10^11m. Find the earth's acceleration towards the sun.


1
Expert's answer
2021-08-03T11:41:11-0400

"v = \frac{s}{T}"

"s = 2\pi r"

"T = 1 \text{ year }\approx 365*24*3600s"

"v= \frac{2\pi*1.5*10^{11}}{365*24*3600}=29.89*10^3\frac{m}{s}"


"a =\frac{v^2}{r}= \frac{29.89^2*10^6}{1.5*10^{11}}=0.00595\frac{m}{s^2}"

"\text{Answer: }a =0.00595\frac{m}{s^2}"